Sociologists say that 80% of married women claim that their husband's mother is the biggest bone of contention in their marriages (sex and money are lower-rated areas of contention). Suppose that eleven married women are having coffee together one morning. Find the following probabilities. (For each answer, enter a number. Round your answers to three decimal places.)

(a)All of them dislike their mother-in-law.

(b)None of them dislike their mother-in-law.

(c) At least nine of them dislike their mother-in-law.

(d) No more than eight of them dislike their mother-in-law.
Answer #1

8- 8080 (1 رواه 0 4 = ال = PC X که از عرع °رو ۵۰ 0 65 3 وہ وہ b) = (م2 اور و.هه ری .قليلا -ها Xلاه . و 1- C P( 29 ) = 2( X =
Answer #2
Sociologists say that 90% of married women claim that their husband’s mother is the biggest bone of contention in their marriages (sex and money are lower-rated areas of contention). (See the source in Problem 18.) Suppose that six married women are having coffee together one morning